OpenStudy (anonymous):

use the following to complete the problems d/dx(sin^(-1)u=(1-u^2)^-1/2(du/dx) d/dx(tan^(-1)u=(1+u)^-1(du/dx) A)y=sin^-1(e^x) B)y=sin^-1[e^(sin^(-1)x] c) y=[tan^-1(x)](1+x^2) I'm not sure what to do at all

OpenStudy (anonymous):

chain rule for each

OpenStudy (anonymous):

so for example, the first one write down the derivative of arcsine, only replace \(x\) by \(e^x\) then multiply your answer by the derivative of \(e^x\) which is \(e^x\)

OpenStudy (anonymous):

first one you get \[\frac{1}{\sqrt{1-(e^x)^2}}\times e^x=\frac{e^x}{\sqrt{1-e^{2x}}}\]

OpenStudy (anonymous):

thank you for your help